OpenStudy (johnjules2000):

Solving multi-step equations: must show work 1/6 (y+42) -15= -3

OpenStudy (a1234):

1/6 (y+42) -15= -3 1/6y + 7 - 15 = -3 1/6y - 8 = -3 1/6y = 5 y = 5/(1/6) y = 30

OpenStudy (johnjules2000):

thanks but does y= 5 1/6 or just 5?

OpenStudy (a1234):

y = 30.

OpenStudy (johnjules2000):

ok thank you

OpenStudy (a1234):

You're welcome.

OpenStudy (johnjules2000):

will you help e with this one ? its the same type of thing 7-y+5y=9

OpenStudy (a1234):

7 - y + 5y = 9 7 + 4y = 9 4y = 2 y = 1/2

OpenStudy (johnjules2000):

Ok thanks again